OpenStudy (anonymous):

Which of the following is the solution to the equation 25^(z + 2) = 125 ?

OpenStudy (lgbasallote):

25 = 5^2 125 = 5^3 so to substitute... 5^2(z+2) = 5^3 2(z+2) = 3 2z + 4 = 3 2z = -1 z = -1/2 got it?

OpenStudy (anonymous):

Thank you!!!!

OpenStudy (anonymous):

You can also, equivalently, take \(\log_{5}\) of both sides. It results in the equation \(2(z+2)=3\).